Two bus A&B travel along the road in same direction from the same starting place. bus A maintains a speed of 120km/h & car B 150km/h but B starts 1hour later how many hours will take for B to over take A?

Grade:12th pass

1 Answers

Saswat Pani
6 years ago
Let the distance when the cars will meet be Sthe time taken by car a is s/120The time taken by car b is s/150And the diff in time is 1 hrSo s/120 - s/150 =1(5s-4s)/600 =1S = 600Time taken by car b= 600/150 = 4hrs